OpenStudy (anonymous):

"the quotient of j and 8" can be expressed as: A. j/8 B. 8j C. j - 8 D. j + 8

OpenStudy (solomonzelman):

quotient is the result of the division, so is I say \[8 \div 2=4\]the quotient is 4 now, \[j \div 8\] since j is an infinity (unknown variable) therefore you can only say that the result is going to be j/8. Which one is it?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

@TheMathDude!

OpenStudy (solomonzelman):

yes, it's A.

OpenStudy (anonymous):

okay i have one more

OpenStudy (anonymous):

Bob and his best friend Bill have the same birthday, but Bob is 3 years older than Bill. Let the variable x represent Bob’s age and y represent Bill’s age. Which equation models the relationship between Bill’s age and Bob’s age? A. y = x/3 B. y = x + 3 C. x = y + 3 D. y = 3x

OpenStudy (texaschic101):

x = y + 3

OpenStudy (texaschic101):

or it could be y = x - 3, but that is not an answer choice

OpenStudy (anonymous):

thanks i need help with one more

OpenStudy (texaschic101):

lets see it

OpenStudy (anonymous):

What is the simplified form of the expression?

OpenStudy (texaschic101):

I believe it is -14

OpenStudy (anonymous):

ok thanks :)

OpenStudy (texaschic101):

sure thing :)